OpenStudy (anonymous):

Consider the line 8x+6y=−5 . What is the slope of a line perpendicular to this line? What is the slope of a line parallel to this line?

OpenStudy (freckles):

Can you tell me the slope of 8x+6y=-5?

OpenStudy (freckles):

If you need help doing that, try rearranging the equation so you have y=mx+b This is called slope-intercept form because it tells us the slope, m and the y-intercept (0,b)

OpenStudy (freckles):

So once you have it in y=mx+b form you know any perpendicular line will have slope -1/m and any parallel line will have slope m